A pion moving with speed 0.6c emits a muon in the same direction. Speed of muon in the rest frame of pion is 

0.4c. calculate the speed of muon in lab frame.


1
Expert's answer
2021-04-08T07:34:03-0400

Explanations & Calculations


  • Since the speeds described here are relativistic, simple velocity addition cannot be employed to calculate the relative speed instead relativistic addition which is given by the equation "\small V_{relative}=\large\frac{U+v^1}{1+\frac{Uv^1}{C^2}}" should be used.


  • Then, how a person on earth sees the motion of the muon is

"\qquad\qquad
\begin{aligned}
\small V&=\small \frac{0.6C+0.4C}{1+\large\frac{0.6C\times0.4C}{C^2}}\\
&=\small \frac{1.0C}{1+0.24}\\
&=\small \bold{0.806C}
\end{aligned}"